You measure 24 textbooks' weights, and find they have a mean weight of 47 ounces. Assume the population standard deviation is 3.6 ounces. Based on this, construct a 95% confidence interval for the true population mean textbook weight. Give your answers as decimals, to two places
Answer #1

Solution

Given that,

= 47

= 3.6

n = 24

A ) At 95% confidence level the z is ,

  = 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.960

Margin of error = E = Z/2* (/n)

= 1.960 * (3.6 / 24 )

= 1.44

At 95% confidence interval estimate of the population mean is,

- E < < + E

47 - 1.44 < < 47 + 1.44

45.56< < 48.44

(45.56 , 48.44 )